OpenStudy (fire212):

Solve by using tables. Give each answer to at most two decimal places. -2x^2-4x=-8x

OpenStudy (xlpoisonpotato):

i would subtract 4x from each side

OpenStudy (xlpoisonpotato):

no wait add

OpenStudy (xlpoisonpotato):

to get -2x^2 = -4x

OpenStudy (xlpoisonpotato):

then divide by -2 to get x^2=2x

OpenStudy (xlpoisonpotato):

then divide by 2 to get 1/2x^2=x